OpenStudy (anonymous):

A sample of 140 firecrackers contained 4 duds. How many duds would you expect in a sample of 2380 firecrackers?

OpenStudy (anonymous):

let x=duds make a proportion 4/140=x/2380 cross multiple 140x=9520 solve for x 9520/140=x x=68 answer 68 duds